# Correlated Proxies: A New Definition and Improved Mitigation for Reward Hacking

### v2

@@ 7 Conclusion
We have introduced a new definition for reward hacking based on correlation between a proxy reward function and the unknown true reward that breaks down when optimizing the proxy. Furthermore, we leveraged this definition to show theoretically and empirically that $\chi^{2}$ occupancy measure regularization can effectively prevent reward hacking. Our results have implications for settings, like RLHF, where RL is used to optimize complex, hard-to-specify objectives. We suggest that the heuristic KL penalty used currently should be replaced by a more principled form of regularization. While OM and AD regularization are equivalent for today’s formulation of RLHF as a contextual bandit, they will no longer remain so as LLM-based agents are optimized over multi-turn conversations or with tool use (Wang et al., 2023; Abdulhai et al., 2023; Shani et al., 2024). Thus, our results provide a principled path to continuing to ensure the safety of increasingly powerful AI systems.
@@ Acknowledgments
We would like to thank Sam Toyer, Micah Carroll, Katie Kang, Dibya Ghosh, and Cam Allen for helpful discussions and feedback on drafts.
@@
Appendix A Proofs
@@
A.2 Understanding the Lower Bound in Theorem 5.1
@@
$L(\pi)=\frac{1}{r}\left(\frac{J(\pi,\tilde{R})-J({\pi_{\text{base}}},\tilde{R})}{\sigma_{\tilde{R}}}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}\right)$

### v3

@@ 7 Conclusion
We have introduced a new definition for reward hacking based on the correlation between a proxy reward function and the unknown true reward that breaks down when optimizing the proxy. Furthermore, we leveraged this definition to show theoretically and empirically that $\chi^{2}$ occupancy measure regularization can more effectively prevent reward hacking than current approaches. Our results have implications for settings, like RLHF, where RL is used to optimize complex, hard-to-specify objectives. We suggest that the heuristic KL penalty used currently should be replaced by a more principled form of regularization. While OM and AD regularization are equivalent for today’s formulation of RLHF as a contextual bandit, they will likely diverge in the near future when LLM-based agents are optimized over multi-turn conversations or with tool use (Wang et al., 2023b; Abdulhai et al., 2023; Shani et al., 2024). Thus, our results provide a principled path to continuing to ensure the safety of increasingly powerful AI systems.
@@ Acknowledgments
We would like to thank Cam Allen, Micah Carroll, Dibya Ghosh, Katie Kang, and Sam Toyer for helpful discussions and feedback on drafts.
@@
Appendix A Proofs and Additional Theoretical Results
@@
A.1.1 Near-Optimal Base Policies
While Theorem 5.1 shows that optimizing the proxy reward with regularization can improve on the base policy’s reward, it does not guarantee that the learned policy will be near-optimal. However, a simple corollary shows that if the base policy is near-optimal, then the learned policy will also be near-optimal.
Corollary A.1.
Suppose that $\tilde{R}$ is an $r$-correlated-proxy for the true reward function $R$, and let $\sigma_{\tilde{R}}$ and $\sigma_{R}$ be defined as in Definition 4.1. Furthermore, suppose that the base policy $\pi_{\text{base}}$} is near-optimal:
$J({\pi_{\text{base}}},R)\geq\max_{\pi^{*}}J(\pi^{*},R)-\epsilon\sigma_{R},$
for some $\epsilon>0$.
Then for any policy $\pi$ such that $\mu_{\pi}\ll\mu_{{\pi_{\text{base}}}}$, we have
$\frac{\max_{\pi^{*}}J(\pi^{*},R)-J(\pi,R)}{\sigma_{R}}\leq\epsilon-\frac{1}{r}\left(\frac{J(\pi,\tilde{R})-J({\pi_{\text{base}}},\tilde{R})}{\sigma_{\tilde{R}}}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}\right).$ (10)
This result bounds the suboptimality gap—how close the learned policy is to optimal—in terms of the suboptimality gap of the base policy and the increase in the regularized proxy reward. The proof is straightforward and follows from Theorem 5.1.
A.1.2 Understanding the Lower Bound in Theorem 5.1
@@ A.1.1 Near-Optimal Base Policies
$L(\pi)=\frac{1}{r}\left(\frac{J(\pi,\tilde{R})-J({\pi_{\text{base}}},\tilde{R})}{\sigma_{\tilde{R}}}-\sqrt{(1-r^{2})\chi^{2}\left(\mu_{\pi}\|\mu_{\pi_{\text{base}}}\right)}\right)$ (11)